OpenStudy (anonymous):

Factor: x^4−16 How do you do them when they are to the 4th power?

OpenStudy (anonymous):

a hint: x^4 = (x^2)^2

OpenStudy (anonymous):

Okay i got that part, but what do you do t the-16? would you do 4^2?

OpenStudy (anonymous):

anything like this: \[x^2 - y^2\] can be factored to this: \[(x+y)(x-y)\] ill show you: \[(x+y)(x-y) = x(x-y) + y(x-y) = x^2 -xy +xy - y^2 = x^2 - y^2\]

OpenStudy (anonymous):

in this case we have \[(x^2)^2 - (4)^2 \] so we factor it to \[(x^2- 4)(x^2 + 4)\] and we can factor again

OpenStudy (anonymous):

Alright that make sense, thanks again!

OpenStudy (anonymous):

no problem :D